Operations Report – January 20th


Crew 220 Operations Report


Submitting Report: Alyx Skywalker, Crew Scientist (Engineer unavailable to report)

Non-nominal Systems: Generator

Generator: Inoperable

Hours Run: 0

Generator running times (Radio Outpost for info) Start Time (HH:MM) = N/A

End Time (HH:MM) = N/A

Total Run time (Hrs) = 0

Solar—SOC %: 0

Diesel Reading: 49%

Station Propane Reading: 25%

Ethanol Free Gasoline: N/A

Water (Loft Tank): 34 gallons in tank

Water Meter: 148,001.8

Water (Static Tank): 487

Static to Loft Pump used – No

Water in GreenHab—245.5 gallons (6.5 gallons for plants, 8 gallons taken to Hab in state of Emergency)

Water in Science Dome: N/A

Toilet Tank Emptied: No


Spirit Rover Used: Yes

Hours: 127.3

Beginning Charge: 100%

Ending Charge: 84%

Currently Charging: Yes

Curiosity Rover Used: Yes

Hours: 133.0

Beginning Charge: 100%

Ending Charge: 87%

Currently Charging: Yes


Spirit Rover Used: Yes

Hours: 127.7

Beginning Charge: 100%

Ending Charge: 78%

Currently Charging: Yes

Curiosity Rover Used: Yes


Beginning Charge: 100%

Ending Charge: 72%

Currently Charging: Yes

Notes on Rovers: None

ATV’s Used: None

HabCar Used and why, where? Did not use today.

General Notes and Comments: None

Summary of Internet: 300mb

Summary of Suits and Radios: Suits are in good working conditions. During our EVAs many of our Radios died despite being fully charged. Currently all the yellow radios work and only 2 blue ones are working. We are charging all radios to test them once charged. Additionally, we only have 2 headsets that worked during our EVAs today.

Summary of Hab Operations: We have been operating on a state of emergency power due to the generator not being functional. We have used minimal lighting and are currently on flashlight power. We have not used Hab power for charging devices today although we did charge the spacesuits, radios and rovers. We did not fill the Loft tank today and have been using as little water that is driven from a pump as possible today to conserve what power we had.

Summary of GreenHab Operations: See GreenHab report

Summary of ScienceDome Operations: No operations today

Summary of RAM Operations: No operations today

Summary of any observatory issues: No issues

Summary of Health and Safety issues: See Health and Safety Report

Questions, concerns and requests to Mission Support: We just request to be kept appraised of the generator situation, especially to be notified when the state of emergency is over or any updates as to the duration of the state of emergency so that we can take appropriate steps.

Journalist Report – January 20th


Crew 220 MDRS Journalist Report 20Jan2020

Author: Connie Delisle, Crew Journalist

Sol 2

Title: Off to the (Mars) Races!

The morning arrived early. The night preceding ended with a flurry of activity in preparing reports, sorting through mission logistics and delineating crew roles. The challenges of ICE Environments are becoming evident, and the crew is already rising to the occasion.

During the morning brief, the MAU and MDRS teams conducted a joint briefing on the proposed joint mission EVAs and mission ops/logistics.

The excitement of this first mission captivated all and many heart-felt shares were offered in wrapping up the meeting. The MAU crew reported making further progress towards station operational efficiency, elated at the challenges and projects ahead.

The morning’s EVA began just before 10:00 am. The MDRS team leader and 2 crew members suited up and walked to the MAU station to join their counterparts. Together, they commenced to shuttle members to the rendezvous point (Marble Ritual). This was followed by a walk around the area prior to returning to their respective stations. This same procedure was repeated for the afternoon EVA which began at 13:30 with the crew members who had stayed behind in the morning. The key difference between EVAs was that the morning’s mission was truncated due to the later start time. Both missions were completed within the set time limit and met mission objectives. Crews are now familiar with the process of donning, doffing, use of the rovers and communications. Follow on EVAs are expected to reinforce familiarity and improve communications consistency.

See photos (separately sent)

Connie Delisle, PhD, MDRS Journalist Crew 220

Copyright © The Mars Society. All rights reserved. | Main Site